# Bibliography.bib

### v1

@@ Abstract.
$\lambda_{1}(-\gamma\Delta+\operatorname{Ric})\geqslant n-1$
for some $0\leqslant\gamma\leqslant\frac{n-1}{n-2}$, then $\operatorname{vol}(M)\leqslant\operatorname{vol}(\mathbb{S}^{n})$, and $\pi_{1}(M)$ is finite. Moreover, the bound on $\gamma$ is sharp for this result to hold. A generalization of the Bonnet–Myers theorem is also shown under the same spectral condition. The proofs involve the use of a new unequally weighted isoperimetric problem and unequally warped $\mu$-bubbles. As an application, in dimensions $3\leqslant n\leqslant 5$, we infer sharp results on the isoperimetric structure at infinity of complete manifolds with nonnegative Ricci curvature and uniformly positive biRicci curvature.
@@ 1.1. Statement of the results
The aim of this paper is to show a spectral generalization of both the classical Bishop–Gromov volume comparison theorem and Bonnet-Myers theorem, and derive some consequences. We stress that in this paper all manifolds $M$ are without boundary.
@@
    Moreover, if equality holds in (3), then every function $u$ as in (1) is constant, and $\widetilde{M}$ is isometric to the round sphere of radius $\lambda^{-1/2}$;
We stress that the range $0\leqslant\gamma\leqslant\frac{n-1}{n-2}$ is sharp, in the sense that for $\gamma>\frac{n-1}{n-2}$ none of the results in Theorem 1 continue to hold. See Remark 3 for a counterexample when $M$ is topologically $\mathbb{S}^{1}\times\mathbb{S}^{n-1}$. In the shorter range $0\leqslant\gamma<\frac{4}{n-1}$ when $n>3$, or in the same range $0\leqslant\gamma\leqslant 2$ when $n=3$, the compactness assumption of $M$ is not needed since it is a consequence of the spectral condition (1).
@@
Let us emphasize the difference between Theorem 1 and Corollary 1: when $n>3$ and for the values $\gamma\in\big{[}\frac{4}{n-1},\frac{n-1}{n-2}\big{)}$, the manifold $M$ satisfying the main spectral condition (1) may not be compact (counterexamples include, for example, hyperbolic spaces). However, assuming that $M$ is compact, we have the volume and diameter bound, and the finiteness of $\pi_{1}(M)$, as a consequence of Theorem 1.
The following conclusion can also be drawn from the proof of Theorem 1: if the manifold $M$ in Theorem 1 is assumed to be only complete, but the function $u$ in (1) is uniformly bounded below and above, then $M$ must be compact as well, see Lemma 1. Indeed, for the known noncompact examples of manifolds satisfying (1) with $\gamma\in\big{[}\frac{4}{n-1},\frac{n-1}{n-2}\big{)}$, the function $u$ either decays at infinity (e.g., hyperbolic spaces) or grows unboundedly at infinity (e.g., the examples in [KaiIntermediate, Subsection 3.3]).
@@
  • For the case $\gamma=0$: the volume bound reduces to the Bishop–Gromov volume comparison theorem, and, in this case, our proof reduces to the isoperimetric comparison argument due to Bray [BraySingular, BrayPhD]. The diameter bound reduces to the Bonnet–Myers theorem, for which our proof provides (up to our knowledge) a new alternative $\mu$-bubble approach.
@@
  • Item 1 becomes $\operatorname{diam}(\widetilde{M})\leqslant\pi/\sqrt{\lambda}$ in dimension $n=3$, thus being a strict spectral generalization of the Bonnet-Myers theorem. The same diameter bound when $n=3$ has appeared with a different proof in [ShenYe, Corollary 1].
@@
  • In the shorter range $0\leqslant\gamma\leqslant 1/(n-2)$, the conclusion about $\pi_{1}(M)$ being finite in Theorem 1 has appeared in [CarronRose, Theorem A(ii)]. The proof given here is different, and the result in Theorem 1 shows that that threshold $0\leqslant\gamma\leqslant 1/(n-2)$ was not sharp, as already guessed by the authors in [CarronRose]. The example in Remark 3 shows that, in Theorem 1, $0\leqslant\gamma\leqslant\frac{n-1}{n-2}$ is the sharp range to get that $\pi_{1}(M)$ is finite.
@@
  • Diameter bounds in presence of spectral conditions as the one in (1) have appeared in [ShenYe, Corollary 1], and [CarronRose, Theorem A(i)].
Remark 1 (Smoothing the eigenfunctions).
Observe that in Theorem 1, and Corollary 1 the function $u$ is assumed to be smooth. On the other hand, calling
$L:=-\gamma\Delta+\mathrm{Ric}-(n-1)\lambda,$
a natural condition that one usually encounters in geometric problems is the nonnegativity of the Dirichlet energy
$\int_{M}\varphi\cdot L\varphi\geqslant 0,\qquad\forall\,\varphi\in C^{\infty}_{c}(M).$ (5)
Since $\operatorname{Ric}$ is only a continuous function, the first eigenfunction of $L$ need not be smooth.
This can be resolved as follows. By regularizing the function $\mathrm{Ric}$, for each $0<\varepsilon<\lambda$ there exists $V_{\varepsilon}\in C^{\infty}(M)$ such that $\|V_{\varepsilon}-\mathrm{Ric}\|_{L^{\infty}(M)}\leqslant\varepsilon(n-1)$. Hence, denoting $L_{\varepsilon}:=-\gamma\Delta+V_{\varepsilon}-(n-1)(\lambda-\varepsilon)$, we have, exploiting (5),
$\int_{M}\varphi\cdot L_{\varepsilon}\varphi\geqslant 0,\qquad\forall\varphi\in C^{\infty}_{c}(M).$
By classical results on elliptic equations (see e.g., [FischerColbrieSchoenCPAM, Theorem 1]), there is a positive function $u_{\varepsilon}\in C^{\infty}(M)$ such that
$\gamma\Delta u_{\varepsilon}=V_{\varepsilon}u_{\varepsilon}-(n-1)(\lambda-\varepsilon)u_{\varepsilon}\leqslant u_{\varepsilon}\mathrm{Ric}-(n-1)\Big{[}\lambda-\varepsilon-\frac{\varepsilon}{n-1}\Big{]}u_{\varepsilon}.$ (6)
Thus, by Theorem 1 and with $\varepsilon\to 0$, we conclude that in the range $0\leqslant\gamma\leqslant\frac{n-1}{n-2}$ it holds:
If $M$ is compact and (5) holds, then $\operatorname{vol}(\widetilde{M})\leqslant\lambda^{-n/2}\operatorname{vol}(\mathbb{S}^{n})$, and $\pi_{1}(M)$ is finite. (7)
The same consequence can be derived from (5) in the range $0\leqslant\gamma<\frac{4}{n-1}$ when $n>3$, without assuming $M$ compact, by using Corollary 1 joint with (6).
Observe that for example the condition in (5) is verified in presence of (small) wells of negative curvature: see, e.g., [ElworthySteven, Proposition 2.1]. ∎
Theorem 1 has consequences on the isoperimetric structure of spaces with nonnegative Ricci curvature and uniformly positive biRicci curvature. Let us recall that, given orthonormal vectors $u,v\in T_{p}M$,
@@
The biRicci curvature was first introduced and studied by Shen–Ye [ShenYeDuke]. There it was shown that in ambient dimensions up to 5, stable minimal hypersurfaces in manifolds with $\mathrm{biRic}\geqslant n-2$ enjoy a uniform diameter bound. This was motivated by and generalizes the 3-dimensional result of Schoen–Yau [SchoenYauBlackHoles]. The underlying idea of [ShenYeDuke] is that the stable minimal surface in question satisfies the spectral control (5) with $\gamma=\lambda=1$. Recently, the study of manifolds with biRicci (and more general intermediate-Ricci) curvature lower bounds has found applications in the solution of the stable Bernstein problem [CLMS], and in the study of a generalization of Geroch conjecture [BrendleJohneHirsch].
@@
where $\partial^{*}E$ denotes the reduced boundary of $E$.
@@
Let $(M^{n},g)$ be a smooth complete noncompact Riemannian manifold with $3\leqslant n\leqslant 5$. Assume that $\mathrm{Ric}\geqslant 0$, and $\mathrm{biRic}\geqslant n-2$ outside a compact set. Suppose further that $M$ has one end. [^1] If not, then $M=\mathbb{R}\times X$, with $\mathrm{Ric}_{X}\geqslant n-2$. Thus $X$ is compact, and $\operatorname{vol}(X)\leqslant\operatorname{vol}(\mathbb{S}^{n-1})$. Hence $M$ has linear volume growth, and $I_{M}(v)\leqslant 2\operatorname{vol}(X)$, with equality for all $v$ large enough. Then the following hold.
@@
    Moreover, equality holds for some $v>0$ in (9), if and only if there is a bounded open set $\Omega\subset M$ such that $M\setminus\Omega$ is isometric to $\mathbb{S}^{n-1}\times[0,+\infty)$.
@@
  • When $n=3$, observe that $\mathrm{biRic}=\mathrm{Scal}/2$. Thus Theorem 2 extends to the noncompact case a former result of Eichmair [EichmairIsop].
@@
    In particular, when $n=3$, Item 1 of Theorem 2 has been obtained in [MunteanuWang22, Theorem 5.6], and generalized (with a different proof) in [ChodoshLiStryker, Theorem 1.1]. See also the recent [WangRecent] for an effective version.
    On the contrary, Item 2 of Theorem 2, which can be interpreted as the sharp isoperimetric control at infinity on the space, is new even in dimension $n=3$. Up to the authors’ knowledge, all the results in Theorem 2 are new when $n\in\{4,5\}$. For a question related to dimension $n\geqslant 6$, see 1.
@@ 1.2. Strategy of the proof
Finally, the proof of Theorem 2 exploits Theorem 1 and the results presented in [ChodoshLiStryker, XingyuZhu, KaiIntermediate]. Roughly speaking, within the range of dimension $3\leqslant n\leqslant 5$, the assumption $\mathrm{biRic}\geqslant n-2$ implies that for every $\varepsilon>0$, one can find a stable $\mu$-bubble $\Omega$ such that
$B_{\varepsilon^{-1}}(o)\subset\Omega,\qquad-\Delta_{\partial\Omega}+\mathrm{Ric}_{\partial\Omega}-(n-2-\varepsilon)\geqslant 0.$ (13)
@@
Let us finally stress one more time that Item 2 of Theorem 1 generalizes to arbitrary dimensions [CLMS, Theorem 5.1], which has been the third part in a three-step-strategy to solve the stable Bernstein problem in $\mathbb{R}^{5}$. Hence the result in Item 2 of Theorem 1 could pave the way to settle the stable Bernstein problem in the remaining dimensions $\mathbb{R}^{6}$, $\mathbb{R}^{7}$.
@@ 1.3. Organization of the paper
In Section 2 we prove Item 1 of Theorem 1. The main tool is Lemma 1, which uses unequally warped $\mu$-bubbles. In Section 3 we prove Item 2 of Theorem 1. First, in Lemma 2 we derive the viscosity ODE (12) for the isoperimetric profile (11). Then the main result is proved by ODE comparison (see Lemma 4). We also discuss the possibility of direct isoperimetry comparison on $\widetilde{M}$ (see Remark 2), and the supercritical case $\gamma>\frac{n-1}{n-2}$ (see Remark 3).
Then in Section 4 we prove Theorem 2. We first produce large sets with small boundary areas in Lemma 5, and then we prove the main ancillary Lemma 6, from which the proof of Theorem 2 follows.
@@ 1.5. Acknowledgments
G.A. is grateful to the Courant Institute for financial support. K.X. would like to thank the hospitality of the Courant Institute, where part of this work was done during his visit. The authors are grateful to Chao Li for interesting discussions and suggestions on the topic of the paper. G.A. is grateful to Marco Pozzetta for useful discussions around the topic of this paper. The authors discussed Theorem 2, which motivated the present paper, when they were participating in the workshop “Recent Advances in Comparison Geometry” in Hangzhou. They thank the Banff International Research Station and the Institute of Advanced Studies in Mathematics for this enriching opportunity.

### v2

@@ Abstract.
$\lambda_{1}\left(-\frac{n-1}{n-2}\Delta+\operatorname{Ric}\right)\geqslant n-1,$
then $\operatorname{vol}(M)\leqslant\operatorname{vol}(\mathbb{S}^{n})$, and $\pi_{1}(M)$ is finite. Moreover, the constant $\frac{n-1}{n-2}$ cannot be improved, and if $\operatorname{vol}(M)=\operatorname{vol}(\mathbb{S}^{n})$ holds, then $M\cong\mathbb{S}^{n}$. A sharp generalization of the Bonnet–Myers theorem is also shown under the same spectral condition.
The proofs involve the use of a new unequally weighted isoperimetric problem, and unequally warped $\mu$-bubbles. As an application, in dimensions $3\leqslant n\leqslant 5$, we infer sharp results on the isoperimetric structure at infinity of complete manifolds with nonnegative Ricci curvature and uniformly positive spectral biRicci curvature.
@@ 1.1. Statement of the results
The aim of this paper is to show a spectral generalization of both the classical Bishop–Gromov volume comparison theorem and Bonnet–Myers theorem, and derive some consequences. We stress that in this paper all manifolds $M$ are without boundary.
@@
    Moreover, if equality holds in (3), then every function $u$ as in (1) is constant, and $\widetilde{M}$ is isometric to the round sphere of radius $\lambda^{-1/2}$.
In the shorter range $0\leqslant\gamma<\frac{4}{n-1}$ when $n>3$, or in the same range $0\leqslant\gamma\leqslant 2$ when $n=3$, the compactness assumption of $M$ in Theorem 1 is not needed since it is implied by the spectral condition (1). Moreover $\mathrm{diam}(M)$ is bounded from above by a constant only depending on $n,\gamma$. The latter assertions are a consequence of the works [ShenYe, KaiIntermediate, CatinoRoncoroni]. So we have the following result.
@@
Remark 1 (Smoothing the eigenfunctions).
Observe that in Theorem 1, and Corollary 1 the function $u$ is assumed to be smooth. On the other hand, a natural condition that one usually encounters is the positivity of the Dirichlet energy
$\int_{M}\gamma|\nabla\varphi|^{2}+\operatorname{Ric}\varphi^{2}\geqslant(n-1)\lambda\int_{M}\varphi^{2},\qquad\forall\,\varphi\in C^{\infty}_{c}(M),$ (5)
for some $\gamma\geqslant 0$, $\lambda>0$. Since $\operatorname{Ric}$ is only a continuous function, the first eigenfunction of $-\gamma\Delta+\mathrm{Ric}$ need not be smooth. Nevertheless, we remark that all the conclusions of Theorem 1 and Corollary 1 remain true under the (a prior weaker) condition (5). This is done by smoothing the first eigenfunction of $-\gamma\Delta+\mathrm{Ric}$, see Remark 5 for the proof.
We remark that, for example, the condition (5) is verified in the presence of small negative curvature wells, or under a Kato condition on $\mathrm{Ric}$: see [ElworthySteven, Proposition 1.2], and [CarronRose], respectively.
Let us comment more thoroughly on the range of $\gamma$.
  1. (i)
    The range $0\leqslant\gamma\leqslant\frac{n-1}{n-2}$ in Theorem 1 is sharp, in the sense that for $\gamma>\frac{n-1}{n-2}$ none of the results there continue to hold. See Remark 4 for a counterexample when $M$ is topologically $\mathbb{S}^{1}\times\mathbb{S}^{n-1}$.
  2. (ii)
    When $n>3$, the range $0\leqslant\gamma<\frac{4}{n-1}$ is optimal to get a universal (i.e., only depending on $n,\gamma$) diameter bound from (1). In fact, in every dimension $n>3$, and for every $\gamma\in\big{(}\frac{4}{n-1},\frac{n-1}{n-2}\big{]}$, there is a compact $M^{n}$ that satisfies $\lambda_{1}(-\gamma\Delta+\operatorname{Ric})\geqslant(n-1)$, but has arbitrarily large diameter. See Remark 2 for the construction.
  3. (iii)
    Let us emphasize the difference between Theorem 1 and Corollary 1: when $n>3$ and for the values $\gamma\in\big{[}\frac{4}{n-1},\frac{n-1}{n-2}\big{]}$, the manifold $M$ satisfying the main spectral condition (1) may not be compact (counterexamples include, for example, hyperbolic spaces). However, assuming that $M$ is compact, we have the volume and diameter bound, and the finiteness of $\pi_{1}(M)$, as a consequence of Theorem 1.
    On the other hand, the following conclusion can be drawn from the proof of Theorem 1: if the manifold $M$ in Theorem 1 is assumed to be only complete, but the function $u$ in (1) is uniformly bounded below and above, then $M$ must be compact as well, see Lemma 1. Indeed, for the known noncompact examples of manifolds satisfying (1) with $n>3$, and $\gamma\in\big{[}\frac{4}{n-1},\frac{n-1}{n-2}\big{]}$, the function $u$ either decays at infinity (e.g., hyperbolic spaces) or grows unboundedly at infinity (e.g., the examples in [KaiIntermediate, Subsection 3.3]).
@@
  1. (i)
    For the case $\gamma=0$: the volume bound in Theorem 1 reduces to the Bishop–Gromov volume comparison theorem, and, in this case, our proof reduces to the isoperimetric comparison argument due to Bray [BraySingular, BrayPhD]. The diameter bound in Theorem 1 reduces to the Bonnet–Myers theorem, for which our proof provides (up to our knowledge) a new alternative $\mu$-bubble approach.
  2. (ii)
@@
  3. (iii)
    Item 1 becomes $\operatorname{diam}(\widetilde{M})\leqslant\pi/\sqrt{\lambda}$ in dimension $n=3$, thus being a strict spectral generalization of the Bonnet–Myers theorem. The same diameter bound when $n=3$ has appeared with a different proof in [ShenYe, Corollary 1].
@@
  5. (v)
    In the shorter range $0\leqslant\gamma\leqslant 1/(n-2)$, the conclusion about $\pi_{1}(M)$ being finite in Theorem 1 has appeared in [CarronRose, Theorem A(ii)]. The proof given here is different, and the result in Theorem 1 shows that that threshold $0\leqslant\gamma\leqslant 1/(n-2)$ was not sharp, as already guessed by the authors in [CarronRose]. As already discussed above, the example in Remark 4 shows that, in Theorem 1, $0\leqslant\gamma\leqslant\frac{n-1}{n-2}$ is the sharp range to get that $\pi_{1}(M)$ is finite.
  6. (vi)
@@
  7. (vii)
    Diameter bounds in presence of spectral conditions as the one in (1) have appeared in [ShenYe, Corollary 1], and [CarronRose, Theorem A(i)]. See [CatinoRoncoroni, Theorem 1.1] for another Bonnet–Myers type theorem involving both spectral and Bakry–Emery Ricci lower bounds.
Theorem 1 has consequences on the isoperimetric structure of spaces with nonnegative Ricci curvature and uniformly positive spectral biRicci curvature. Let us recall that, given orthonormal vectors $u,v\in T_{p}M$,
@@
The biRicci curvature was first introduced and studied by Shen–Ye [ShenYeDuke]. There it was shown that in ambient dimensions up to 5, stable minimal hypersurfaces in manifolds with $\mathrm{biRic}\geqslant n-2$ enjoy a uniform diameter bound. This was motivated by and generalizes the 3-dimensional result of Schoen–Yau [SchoenYauBlackHoles]. The underlying idea of [ShenYeDuke] is that the stable minimal surface in question satisfies the spectral control (5) with $\gamma=1$, and a positive $\lambda>0$. Recently, the study of manifolds with biRicci (and more general intermediate-Ricci) curvature lower bounds has found applications in the solution of the stable Bernstein problem [CLMS], and in the study of a generalization of Geroch conjecture [BrendleJohneHirsch].
@@
where $\partial^{*}E$ denotes the reduced boundary of $E$. Let us denote
$\mathrm{biRic}(x):=\inf_{v,w\in T_{x}M\,\text{orthonormal}}\mathrm{biRic}_{x}(v,w),$
and let us say that, for constants $\gamma\geqslant 0,\lambda>0$, a manifold satisfies $\lambda_{1}(-\gamma\Delta+\operatorname{biRic})\geqslant\lambda$ outside a compact set, if there is a compact set $K\subset M$ such that
$\int\gamma|\nabla\varphi|^{2}+\operatorname{biRic}\varphi^{2}\geqslant\lambda\int\varphi^{2},\qquad\forall\varphi\in C^{\infty}_{c}(M\setminus K).$
@@
Let $(M^{n},g)$ be a smooth complete noncompact Riemannian manifold with $3\leqslant n\leqslant 5$, and let $0\leqslant\gamma<6-n$ (or $0\leqslant\gamma\leqslant 2$ when $n=4$). Assume that $\mathrm{Ric}\geqslant 0$ and $\lambda_{1}(-\gamma\Delta+\mathrm{biRic})\geqslant n-2$ outside a compact set. Assume further that $M$ has one end. [^1] If not, then $M=\mathbb{R}\times X$, with $\lambda_{1}(-\gamma\Delta+\mathrm{Ric}_{X})\geqslant n-2$. Thus $X$ is compact, and $\operatorname{vol}(X)\leqslant\operatorname{vol}(\mathbb{S}^{n-1})$ by Corollary 1. Hence $M$ has linear volume growth, and $I_{M}(v)\leqslant 2\operatorname{vol}(X)$, with equality for all $v$ large enough. Then the following hold.
@@
    Moreover, equality of (7) holds for some $v>0$ if and only if there is a bounded open set $\Omega\subset M$ such that $M\setminus\Omega$ is isometric to $\mathbb{S}^{n-1}\times[0,+\infty)$.
@@
  1. (i)
    When $n=3$, observe that $\mathrm{biRic}=\mathrm{Scal}/2$. Thus, when $n=3$ and $\gamma=0$, Theorem 2 extends to the noncompact case a former result of Eichmair [EichmairIsop].
  2. (ii)
@@
    In particular, when $n=3$ and $\gamma=0$, Item 1 of Theorem 2 has been obtained in [MunteanuWang22, Theorem 5.6], and generalized (with a different proof) in [ChodoshLiStryker, Theorem 1.1]. See also the recent [WangRecent] for an effective version.
    On the contrary, Item 2 of Theorem 2, which can be interpreted as the sharp isoperimetric control at infinity on the space, is new even when $n=3$ and $\gamma=0$. Up to the authors’ knowledge, Item 1 of Theorem 2 is new when $n\in\{4,5\}$, even when $\gamma=0$. For a question related to dimension $n\geqslant 6$, see 1.
  3. (iii)
@@ 1.2. Strategy of the proof
Finally, the proof of Theorem 2 exploits Theorem 1 and results and idea presented in [ChodoshLiStryker, XingyuZhu, KaiIntermediate, CLMS]. Roughly speaking, within the range of dimension $3\leqslant n\leqslant 5$, and when $0\leqslant\gamma<6-n$ (or $0\leqslant\gamma\leqslant 2$ when $n=4$), the assumption $\lambda_{1}(-\gamma\Delta+\mathrm{biRic})\geqslant n-2$ implies that for every $\varepsilon>0$, one can find a stable warped $\mu$-bubble $\Omega$ such that
$B_{\varepsilon^{-1}}(o)\subset\Omega,\qquad-\frac{4}{4-\gamma}\Delta_{\partial\Omega}+\mathrm{Ric}_{\partial\Omega}-(n-2-\varepsilon)\geqslant 0.$ (11)
@@
Added notes. Besides the application discussed in Theorem 2, we expect the results of this paper to be useful to deal with a variety of geometric problems. After this paper was made public, the main result Theorem 1, together with the strategy devised in [CLMS], has been exploited by L. Mazet to settle the stable Bernstein problem in $\mathbb{R}^{6}$ [Mazet]. Theorem 1 could pave the way to study the stable Bernstein problem in the remaining dimension $\mathbb{R}^{7}$. In addition, it would be interesting to understand the implications of our results to the study of the structure of manifolds (or limits of manifolds) with uniform Kato bounds on the Ricci curvature: see, e.g., [CarronMondelloTewodrose].
@@ 1.3. Organization of the paper
In Section 2 we prove Item 1 of Theorem 1. The main tool is Lemma 1, which uses unequally warped $\mu$-bubbles. Then in Remark 2, we discuss the possibility of improving (2) in the range $\gamma\in\left[\frac{4}{n-1},\frac{n-1}{n-2}\right]$. In Section 3 we prove Item 2 of Theorem 1. First, in Lemma 2 we derive the viscosity ODE (10) for the isoperimetric profile (9). Then the main result is proved by ODE comparison (see Lemma 4). We also discuss the possibility of direct isoperimetry comparison on $\widetilde{M}$ (see Remark 3), the supercritical case $\gamma>\frac{n-1}{n-2}$ (see Remark 4), and we prove Remark 1 in Remark 5.
Then in Section 4 we prove Theorem 2. We first produce large sets with small boundary areas by means of Lemma 5, and then we prove the main ancillary Lemma 6, from which the proof of Theorem 2 follows. We also discuss the implications of the results in Theorem 2 to sharp linear volume growth of the manifold at infinity in Remark 7.
@@ 1.5. Acknowledgments
G.A. acknowledges the financial support of the Courant Institute, and the AMS-Simons Travel grant. K.X. would like to thank the hospitality of the Courant Institute, where part of this work was done during his visit. The authors are grateful to Chao Li, Xian-Tao Huang, Marco Pozzetta, and Alberto Roncoroni for interesting discussions and suggestions on the topic of the paper. The authors are grateful to Otis Chodosh for comments that led to the improvement of a previous version of Theorem 2, and to Jintian Zhu for comments that led to Remark 2.
The authors discussed Theorem 2, which motivated the present paper, when they were participating in the workshop “Recent Advances in Comparison Geometry” in Hangzhou. They thank the Banff International Research Station and the Institute of Advanced Studies in Mathematics for this enriching opportunity.